# Evolution with Opponent-Learning Awareness

### v2

@@ 1. Introduction
While some notable achievements do tackle MARL scenarios, these are largely constrained by the high inherent complexity of multi-agent training. Successful MARL applications rarely feature more than a few agents, and in fact often resort to training a single neural network against one or few copies of itself. This powerful learning trick, named “self-play”, was famously used to learn Go at a superhuman level Silver et al. (2016), and competitive human-level team play in a video game Berner et al. (2019). On both occasions, the authors had to spend extreme amounts of compute, even though the number of agents involved was small and despite the computational savings enabled by training a single replicated model (as opposed to independent agents). Furthermore, recent successful MARL achievements typically relied on said “Multi-Agent” versions of single-agent RL algorithms (MADDPG Lowe et al. (2020), MAPPO Yu et al. (2022)…) which fall under a category of MARL techniques broadly referred to as “centralized training” Wong et al. (123). Centralized training consist of making use of privileged global information during training to reduce the fundamental non-stationarity provoked by the evolution of others, for instance by feeding this information to a centralized critic.
@@
The selection and replication heuristics used in ES are inspired from “imitation of the fittest”, a gradient-free revision protocol through which populations of non-learning agents are thought to evolve in nature Apesteguia et al. (2007). When aggregated across many agents, imitation of the fittest yields the Replicator population dynamic Sandholm (2010), arguably the most famous equation in Evolutionary Game Theory (EGT) Weibull (1997). Whereas MARL is often practically constrained to few-agent scenarios, EGT has long studied the evolution of large populations made of independent non-learning agents. Because these agents are non-learning, the revision protocols considered in EGT are extrinsic, similar to the genetically-inspired replicating and pruning performed in ES. However, imitation of success can also be considered as a very simple intrinsic learning protocol, where the agent simply copies the policy of its opponent based on their payoff. In this work, we bring advanced intrinsic MARL protocols to the scale of evolutionary simulations. In particular, we are interested in the evolutionary effects of Opponent-Learning Awareness, an advanced MARL protocol able to take advantage of non-stationary dynamics Foerster et al. (2017). Our contribution is threefold:
@@ 2.1. Evolutionary Game Theory
EGT models evolution as a series of random pairwise interactions, where interactions are typically simple bi-matrix games (i.e., 2-player multi-armed bandits). Agents are sampled from a large population to be randomly paired and evaluated against their drawn opponent. The outcome of this interaction is a payoff for each opponent, whose expectation is called the agent’s fitness against the current population. An agent’s fitness depends on both its policy and the current configuration of the population. The agent’s policy is called its type, which is one of $n$ possible types. Agents with a greater fitness replicate and thus tend to “invade” the population, whereas agents with a lower fitness tend to go extinct. In particular, EGT is interested in evolutionarily stable equilibria, which are configurations of the population where the different types are present in stable proportions under replication dynamics. In evolutionarily stable equilibria, the population configuration is robust to rare mutations, where few agents randomly switch from one type to another. In this paper, we will be studying more complex, learning-based types of evolutionary equilibria in three classic symmetric games: Stag Hunt, Hawk-Dove, and Rock-Paper-Scissors.
@@ 2.1.1. Stag hunt
where rows represent the action chosen by the ego agent (bold) with corresponding payoffs in first position, and columns represent the action chosen by the other agent with corresponding payoffs in second position. The game of Hawk-Dove has two distinct pure strategy Nash equilibria2222-player equilibria where each agent’s best response to the other player’s strategy is to always deterministically select the same action. Nash et al. (1950): (1) both agents always playing Stag, and (2) both agents always playing Hare.
@@ 2.1.2. Hawk-Dove
The game of Hawk-Dove has two pure strategy Nash equilibria: (1) Hawk-Dove and (2) Dove-Hawk. But note that in these equilibria, Doves have a smaller payoff than Hawks. In the context of EGT, this means that when Doves encounter almost only Hawks, they go extinct and Hawks invade. However, when Hawks encounter almost always Hawks, their payoff is even less than Doves encountering Hawks, and thus Hawks move toward extinction as Doves invade. In other words, there are population configurations in which it is not more relevant to be a Hawk than a Dove in terms of fitness, and evolutionary dynamics naturally drive the population there. Examples of such configurations are (1) a population where all agents pick Hawk or Dove uniformly at random (which is the mixed-strategy Nash equilibrium3332-player equilibrium where both agents assign probabilities to each action. of this game), (2) a population formed of half deterministic Hawks and half deterministic Doves, or (3) any arbitrary population whose average behavior is to chose Hawk and Dove uniformly at random (for instance, half agents with policy $(\frac{2}{3},\frac{1}{3})$ and half agents with policy $(\frac{1}{3},\frac{2}{3})$…).
@@ 2.1.4. Population vector
A population whose individual are distributed amongst $n$ types can be represented as a population vector $P\in\mathbb{R}^{n}$:
@@ 2.1.5. Replicator Dynamic
Under the “imitation of the fittest” revision protocol (as well as several other bio-inspired revision protocols), large populations are known to follow a famous population dynamic, called the Replicator Dynamic:
@@ 3. Methods
Whereas these rules are usually thought of in the context of persistent interactions between fixed agent pairs, in the evolutionary setting, they instead get applied after single interactions between random pairs. In other words, learning agents consider that the random opponent they interact with at each evolution step is a representative sample of the population and, for LOLA, of the current direction of its non-stationary dynamics. As hinted in Section 2.1, this yields complex population dynamics that may of may not converge to homogeneous populations, depending on factors such as the game, the learning rule, and the policy architecture.
@@ 3.2. Analytical Policy Gradient
This formulation yields the following analytical PG formulation for symmetric normal-games:
@@ 4.5. Rock-Paper-Scissors
Figure 8 presents the results of our evolutionary simulations in the RPS game. The color-coding follows the same principle as our previous population plots, with the bottom triangle being the policy, and evolution steps being the vertical dimension. Similarly to Figure 6(b), Figure 8(a) shows that populations of naive learners evolve into 3 equally distributed groups of close-to-deterministic agents always playing Rock, Paper and Scissors respectively. The reason why this happens is of interest, and is more clearly understood from Figure 9 in Appendix A.2. In short, this dynamic results from the loss of plasticity modelled by the softmax function [^11] This model is very similar to the “cost of motion” described by Mertikopoulos and Sandholm for the Replicator Dynamic Mertikopoulos and Sandholm (2018).. At the beginning of evolution, naive learners are erratically moving around as they encounter all types of strategies. However, after some time, agents get “trapped” near the border of the policy simplex, where gradients toward the opposite action are near-zero. After a long time, three groups of near-deterministic agents emerge, and a small number of them continuously escape toward the strategy that counters the majority group, which eventually creates a new majority, and so on, yielding a cyclic evolution pattern. In other words, diversity emerges from populations of naive learners in the RPS model. On the other hand, Figure 8(b) tells a different story about populations of LOLA agents, which instead quickly and unanimously converge to the mixed Nash equilibrium of this game.
@@ 4.6. Remark
From our simulation results, it looks like the mean policy averaged over the entire population always converges near a Nash equilibrium of the game, even when the learning rule itself does not converge to this equilibrium in the conventional 2-agent MARL setting (see Figures 5(b) and 7(a)). We believe that this interesting property is however merely a consequence of the uniform random opponent matching scheme that we chose to implement in this paper. This can be seen by considering an extreme opposite scheme, where all pairs would instead interact persistently, as all individual pairs would then converge to the pure Nash equilibrium in the Hawk-Dove game (that is, exactly one deterministic Hawk and one deterministic Dove per pair). This would average to a uniform random policy, which only happens to coincide with the mixed Nash equilibrium in our example because of our particular choice of payoff for the Hawk/Hawk situation. Any slight modification to the cost of fighting would induce a different mixed Nash equilibrium, but would not change the pure Nash equilibrium of this game. In reality, partner selection is more structured Anastassacos et al. (2020) and can lead to different outcomes, which we plan to explore in the future.
@@ A.1. Second-order policy gradients
This enables writing the LOLA second order gradient as:

### v3

@@ 1. Introduction
While some notable achievements do tackle MARL scenarios, these are largely constrained by the high inherent complexity of multi-agent training. Successful MARL applications rarely feature more than a few agents, and in fact often resort to training a single neural network against one or few copies of itself. This powerful learning trick, named “self-play”, was famously used to learn Go at a superhuman level Silver et al. (2016), and competitive human-level team play in a video game Berner et al. (2019). On both occasions, the authors had to spend extreme amounts of compute, even though the number of agents involved was small and despite the computational savings enabled by training a single replicated model (as opposed to independent agents). Furthermore, recent successful MARL achievements typically relied on said “Multi-Agent” versions of single-agent RL algorithms (MADDPG Lowe et al. (2020), MAPPO Yu et al. (2022)…) which fall under a category of MARL techniques broadly referred to as “centralized training” Wong et al. (123). Centralized training consists of making use of privileged global information during training to reduce the fundamental non-stationarity provoked by the evolution of others, for instance by feeding this information to a centralized critic.
@@
The selection and replication heuristics used in ES are inspired by “imitation of the fittest”, a gradient-free revision protocol through which populations of non-learning agents are thought to evolve in nature Apesteguia et al. (2007). When aggregated across many agents, imitation of the fittest yields the Replicator population dynamic Sandholm (2010), arguably the most famous equation in Evolutionary Game Theory (EGT) Weibull (1997). Whereas MARL is often practically constrained to few-agent scenarios, EGT has long studied the evolution of large populations made of independent non-learning agents. Because these agents are non-learning, the revision protocols considered in EGT are extrinsic, similar to the genetically-inspired replicating and pruning performed in ES. However, imitation of success can also be considered as a very simple intrinsic learning protocol, where the agent simply copies the policy of its opponent based on their payoff. In this work, we bring advanced intrinsic MARL protocols to the scale of evolutionary simulations. In particular, we are interested in the evolutionary effects of Opponent-Learning Awareness, an advanced MARL protocol able to take advantage of non-stationary dynamics Foerster et al. (2017). Our contribution is threefold:
@@ 2.1. Evolutionary Game Theory
EGT models evolution as a series of random pairwise interactions, where interactions are typically simple bi-matrix games (i.e., 2-player multi-armed bandits). Agents are sampled from a large population to be randomly paired and evaluated against their drawn opponent. The outcome of this interaction is a payoff for each opponent, whose expectation is called the agent’s fitness against the current population. An agent’s fitness depends on both its policy and the current configuration of the population. The agent’s policy is called its type, which is one of $n$ possible types. Agents with a greater fitness replicate and thus tend to “invade” the population, whereas agents with a lower fitness tend to go “extinct”. In particular, EGT is interested in evolutionarily stable equilibria, which are configurations of the population where the different types are present in stable proportions under replication dynamics. In evolutionarily stable equilibria, the population configuration is robust to rare mutations, where few agents randomly switch from one type to another. In this paper, we will be studying more complex, learning-based types of evolutionary equilibria in three classic symmetric games: Stag Hunt, Hawk-Dove, and Rock-Paper-Scissors.
@@ 2.1.1. Stag hunt
where rows represent the action chosen by the ego agent (bold) with corresponding payoffs in first position, and columns represent the action chosen by the other agent with corresponding payoffs in second position. The game of Stag hunt has two distinct pure strategy Nash equilibria2222-player equilibria where each agent’s best response to the other player’s strategy is to always deterministically select the same action. Nash et al. (1950): (1) both agents always playing Stag, and (2) both agents always playing Hare.
@@ 2.1.2. Hawk-Dove
The game of Hawk-Dove has two pure strategy Nash equilibria: (1) Hawk-Dove and (2) Dove-Hawk. But note that in these equilibria, Doves have a smaller payoff than Hawks. In the context of EGT, this means that when Doves encounter almost only Hawks, they go extinct and Hawks invade. However, when Hawks encounter almost always Hawks, their payoff is even less than Doves encountering Hawks, and thus Hawks move toward extinction as Doves invade. In other words, there are population configurations in which it is not more relevant to be a Hawk than a Dove in terms of fitness, and evolutionary dynamics naturally drive the population there. Examples of such configurations are (1) a population where all agents pick Hawk or Dove uniformly at random (which is the mixed-strategy Nash equilibrium3332-player equilibrium where both agents assign probabilities to each action. of this game), (2) a population formed of half deterministic Hawks and half deterministic Doves, or (3) any arbitrary population whose average behavior is to choose Hawk and Dove uniformly at random (for instance, half agents with policy $(\frac{2}{3},\frac{1}{3})$ and half agents with policy $(\frac{1}{3},\frac{2}{3})$…).
@@ 2.1.4. Population vector
A population whose individuals are distributed amongst $n$ types can be represented as a population vector $P\in\mathbb{R}^{n}$:
@@ 2.1.5. Replicator Dynamic
Under the “imitation of the fittest” revision protocol (as well as several other bio-inspired revision protocols), large populations are known to follow a famous population dynamic over time ($t$), called the Replicator Dynamic:
@@ 3. Methods
Whereas these rules are usually thought of in the context of persistent interactions between fixed agent pairs, in the evolutionary setting, they instead get applied after single interactions between random pairs. In other words, learning agents consider that the random opponent they interact with at each evolution step is a representative sample of the population and, for LOLA, of the current direction of its non-stationary dynamics. As hinted in Section 2.1, this yields complex population dynamics that may or may not converge to homogeneous populations, depending on factors such as the game, the learning rule, and the policy architecture.
@@ 3.2. Analytical Policy Gradient
This formulation yields the following analytical PG formulation for symmetric normal-form games:
@@ 4.5. Rock-Paper-Scissors
Figure 8 presents the results of our evolutionary simulations in the RPS game. The color-coding follows the same principle as our previous population plots, with the bottom triangle being the policy, and evolution steps being the vertical dimension. Similarly to Figure 6(b), Figure 8(a) shows that populations of naive learners evolve into 3 equally distributed groups of close-to-deterministic agents always playing Rock, Paper and Scissors respectively. The reason why this happens is of interest, and is more clearly understood from Figure 9 in Appendix A.2. In short, this dynamic results from the loss of plasticity modeled by the softmax function [^11] This model is very similar to the “cost of motion” described by Mertikopoulos and Sandholm for the Replicator Dynamic Mertikopoulos and Sandholm (2018).. At the beginning of evolution, naive learners are erratically moving around as they encounter all types of strategies. However, after some time, agents get “trapped” near the border of the policy simplex, where gradients toward the opposite action are near-zero. After a long time, three groups of near-deterministic agents emerge, and a small number of them continuously escape toward the strategy that counters the majority group, which eventually creates a new majority, and so on, yielding a cyclic evolution pattern. In other words, diversity emerges from populations of naive learners in the RPS model. On the other hand, Figure 8(b) tells a different story about populations of LOLA agents, which instead quickly and unanimously converge to the mixed Nash equilibrium of this game.
@@ 4.6. Remark
From our simulation results, it looks like the mean policy averaged over the entire population always converges near a Nash equilibrium of the game, even when the learning rule itself does not converge to this equilibrium in the conventional 2-agent MARL setting (see Figures 5(b) and 7(a)). We believe that this interesting property is however merely a consequence of the uniform random opponent matching scheme that we chose to implement in this paper. For instance, let us consider an extreme opposite scheme, where all pairs would instead interact persistently. All individual pairs would then converge to the pure Nash equilibrium in the Hawk-Dove game (that is, exactly one deterministic Hawk and one deterministic Dove per pair). This would average to a uniform random policy, which only happens to coincide with the mixed Nash equilibrium in our example because of our particular choice of payoff for the Hawk/Hawk situation. Any slight modification to the cost of fighting would induce a different mixed Nash equilibrium, but would not change the pure Nash equilibrium of this game. In reality, partner selection is more structured Anastassacos et al. (2020) and can lead to different outcomes, which we plan to explore in the future.
@@ A.1. Second-order policy gradients
This enables writing the LOLA second-order gradient as: